# DiffTester: Accelerating Unit Test Generation for Diffusion LLMs via Repetitive Pattern

### v1

@@ C.1.3 Hyperparameters
In our experiments with the EB-Sampler method, we adopted the parameters recommended in the original paper Ben-Hamu et al. (2025b). We used confidence as the error proxy function. And we used $\gamma=0.1$.
@@ C.1.4 Benchmark
To evaluate the performance of our method, we experiment on the TestEval benchmark Wang et al. (2025). TestEval is a benchmark for test case generation with LLMs. It includes 210 Python programs from an online programming platform, LeetCode. It has three metrics: overall coverage, targeted line/branch coverage, and targeted path coverage.
@@ C.5 Limitation
A principal limitation of DiffTester is the absence of empirical investigations into the acceleration of UTG on repository-level codebases, which would be of greater practical significance. This omission is primarily attributable to the constrained maximum context lengths supported by existing dLLM architectures. In contrast, repository-scale UTG tasks, such as those exemplified by TestGenEval (Jain et al., 2025), inherently demand the modeling of substantially longer contexts that exceed these architectural limits.
Another limitation arises from the inherent characteristics of certain contemporary code generation models, such as DreamCoder (Xie et al., 2025). Our method was not applied to these models primarily because their generative behavior tends to produce substantial amounts of natural language output, even when explicitly prompted to focus on code generation. This results in a low proportion of code segments, thereby constraining the potential acceleration benefits achievable by our method.

### v2

@@ C.1.3 Hyperparameters
In our experiments with the EB-Sampler method, we adopted the parameters recommended in the original paper ben2025accelerated. We used confidence as the error proxy function. And we used $\gamma=0.1$.
@@ C.1.4 Benchmark
To evaluate the performance of our method, we experiment on the TestEval benchmark testeval. TestEval is a benchmark for test case generation with LLMs. It includes 210 Python programs from an online programming platform, LeetCode. It has three metrics: overall coverage, targeted line/branch coverage, and targeted path coverage.
@@ C.5 Limitation
A principal limitation of DiffTester is the absence of empirical investigations into the acceleration of UTG on repository-level codebases, which would be of greater practical significance. This omission is primarily attributable to the constrained maximum context lengths supported by existing dLLM architectures. In contrast, repository-scale UTG tasks, such as those exemplified by TestGenEval (jain2025testgenevalrealworldunit), inherently demand the modeling of substantially longer contexts that exceed these architectural limits.
Another limitation arises from the inherent characteristics of certain contemporary code generation models, such as DreamCoder (dreamcoder). Our method was not applied to these models primarily because their generative behavior tends to produce substantial amounts of natural language output, even when explicitly prompted to focus on code generation. This results in a low proportion of code segments, thereby constraining the potential acceleration benefits achievable by our method.